Show that( R [x] /<x²+1>) isomorphic To C as fields.
1
Expert's answer
2020-11-25T17:55:57-0500

We give a map "f:\mathbb{R}[X]\longrightarrow \mathbb{C}." "x\mapsto i." In general, "g(x)\mapsto g(i)." Hence "a+bX\mapsto a+ib." Hence surjective. Homomorphism is obvious since "gh(i)=g(i)h(i)" for polynomials. Also "(g+h)(i)=g(i)+h(i)." Now for kernel. "g(x)\mapsto 0\Rightarrow g(i)=0." Now by division algorithm, "g(x)=aX+b+(X^2+1)(q(x))" . Hence "g(i)=0\Rightarrow ai+b=0\Rightarrow a=b=0." Hence

kernel is generated by "X^2+1." Hence by first isomorphism theorem, the result follows.